The diagonal of a rectangular field is 60 metres more than the shorter side. If the longer side is 30 metres more than the shorter side, find the sides of the field.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us say, the shorter side of the rectangle bexm.
Then, larger side of the rectangle = (x+ 30) m

the length of the diagonal is = [ (x + 30)^2 + x^2 ]^(1/2) m
Given, diagonal= x+60

Therefore,

⇒x^2+(x+30)^2=(x+60)^2
⇒x^2+x^2+900+60x=x^2+3600+120x
⇒x^2- 60x- 2700 = 0
⇒x^2-90x+30x-2700=0
⇒x(x- 90) + 30(x-90) = 0
⇒(x- 90)(x+ 30) = 0
⇒x= 90, -30

Thanks